OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3In a-1/2(In b + In c^2) @kropot72

OpenStudy (mertsj):

lna^3-1/2lnb-1/2lnc^2= lna^3-lnsqrtb-lnc=ln(a^3/csqrtb)